Assume that the MPC is .8, calculate the effects of an increase in G of $30B on output, consumption and savings.
1
Expert's answer
2017-06-17T04:14:10-0400
∆Y=∆G*1/(1-MPC)=30*5=150B

150=∆C*5

∆C=30

∆S=∆I=∆Y-∆C-∆G=150-30- 30=90B